OpenStudy (anonymous):

Write the first ten terms of a sequence whose first term is -10 and whose common difference is -2.

OpenStudy (anonymous):

a\[a _{n} = a + d; a _{n} = a+ d(n-1)\] a= -10 d= -2 n = 1 - 10

OpenStudy (jdoe0001):

-10 -10 -2 (-10 -2) -2 (-10 -2 -2 ) -2 ....

OpenStudy (jdoe0001):

do so up to the 10th term